OpenStudy (anonymous):

Solve the equation e(power x) - e(power -x)=6

jhonyy9 (jhonyy9):

e^x -e^(-x) =6 e^x - 1/e^x =6 e^2x -1=6e^x e^2x -6e^x -1 =0 note e^x =y - so than y^2 -6y -1 =0 6 +/- radical(36+4) 2(3+/- radical 10) y_1,_2 = ------------------ = ---------------- = 3+/- radical 10 2 2 - after this you need to resolving 2 cases : 1. when e^x = 3+radical 10 2. when e^x = 3-radical 10

OpenStudy (anonymous):

well the any key is a. In(6+square root 10) b. In(3+square root 10) c. In(1+square root 10) d. Insquare root 10 e. In(2+square root 10

jhonyy9 (jhonyy9):

yes ln(3+/- sqrt 10)

jhonyy9 (jhonyy9):

- so can you understanding all ?

jhonyy9 (jhonyy9):

if yes your secondly exercise you can resolving the same like this successes

OpenStudy (anonymous):

well the any key is a. In(6+square root 10) b. In(3+square root 10) c. In(1+square root 10) d. Insquare root 10 e. In(2+square root 10

OpenStudy (anonymous):

on step number three how did you get e to the power 2x?

jhonyy9 (jhonyy9):

so this is easy looke you need to check the denominator common and after this you need to multiplie the terms without this denominator with what you have got like

jhonyy9 (jhonyy9):

so this is clearly ?

jhonyy9 (jhonyy9):

because you have e^(-x) this sign that this is equal 1/e^x and hence there is one denominator

jhonyy9 (jhonyy9):

now do you understanding ?

OpenStudy (anonymous):

I am still kind of confused?